Mark can row upstream at 5mph and downstream at 11 mph if mark starts rowing upstream until he gets tired and then rows downstream to his starting point how far did he row if the entire trip took 4 hrs

Answer by lwsshak3(11628)

Mark can row upstream at 5mph and downstream at 11 mph if mark starts rowing upstream until he gets tired and then rows downstream to his starting point how far did he row if the entire trip took 4 hrs
***
let x=travel time upstream
4-x=travel time downstream
distance=travel time*speed
5x=11(4-x)
5x=44-11x
16x=44
x=44/16=11/4
4-x=16/4-11/4=5/4
5*11/4=55/4(distance upstream)
11*5/4=55/4(distance downstream)
2*55/4=55/2=27.5
how far did he row? 27.5 miles